Solve the given problems



1. A population consists of the four numbers 1. 2. 4 and 5. List all the possible



samples of size n = 3 which can be drawn with replacement from the population



Find the following



a Population mean



b. Population variance



c. Population standard deviation



d. Mean of the sampling distribution of sample means



e. Variance of the sampling distribution of sample means



f Standard deviation of the sampling distribution of sample means

1
Expert's answer
2022-05-19T06:03:36-0400

We have population values "1,2,4,5" population size "N=4" and sample size "n=3."

Thus, the number of possible samples which can be drawn with replacement is "4^3=64."

a. Population mean:


"\mu=\dfrac{1+2+4+5}{4}=3"


b. Population variance:



"\sigma^2=\dfrac{1}{4}((1-3)^2+(2-3)^2+(4-3)^2+(5-3)^2)""=2.5"



c. Population standard deviation:



"\sigma=\sqrt{\sigma^2}=\sqrt{2.5}\approx1.58114"


d.



"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Sample\ values & Sample\ mean(\bar{X}) \\ \hline
 1,1,1 & 1\\
 \hdashline
 1,1,2 & 4/3\\
 \hdashline
 1,1,4 & 2\\
 \hdashline
 1,1,5 & 7/3\\
 \hdashline
 1,2,1 & 4/3\\
 \hdashline
 1,2,2 & 5/3\\
 \hdashline
 1,2,4 & 7/3\\
 \hdashline
 1,2,5 & 8/3\\
 \hdashline
 1,4,1 & 2\\
 \hdashline
 1,4,2 & 7/3\\
 \hdashline
 1,4,4 & 3\\
 \hdashline
 1,4,5 & 10/3\\
 \hdashline
 1,5,1 & 7/3\\
 \hdashline
 1,5,2 & 8/3\\
 \hdashline
 1,5,4 & 10/3\\
 \hdashline
 1,5,5 & 11/3\\
 \hdashline
 2,1,1 & 4/3\\
 \hdashline
 2,1,2 & 5/3\\
 \hdashline
 2,1,4 & 7/3\\
 \hdashline
 2,1,5 & 8/3\\
 \hdashline
 2,2,1 & 5/3\\
 \hdashline
 2,2,2 & 2\\
 \hdashline
 2,2,4& 8/3\\
 \hdashline
 2,2,5 & 3\\
 \hdashline
 2,4,1 & 7/3\\
 \hdashline
 2,4,2 & 8/3\\
 \hdashline
 2,4,4 & 10/3\\
 \hdashline
 2,4,5 & 11/3\\
 \hdashline
 2,5,1 & 8/3\\
 \hdashline
 2,5,2 & 3\\
 \hdashline
 2,5,4 & 11/3\\
 \hdashline
 2,5,5 & 4\\
 \hdashline
 4,1,1 & 2\\
 \hdashline
 4,1,2 & 7/3\\
 \hdashline
 4,1,4 & 3\\
 \hdashline
 4,1,5 & 10/3\\
 \hdashline
 4,2,1 & 7/3\\
 \hdashline
 4,2,2 & 8/3\\
 \hdashline
 4,2,4 & 10/3\\
 \hdashline
 4,2,5 & 11/3\\
 \hdashline
 4,4,1 & 3\\
 \hdashline
 4,4,2 & 10/3\\
 \hdashline
 4,4,4 & 4\\
 \hdashline
 4,4,5 & 13/3\\
 \hdashline
 4,5,1 & 10/3\\
 \hdashline
 4,5,2 & 11/3\\
 \hdashline
 4,5,4 & 13/3\\
 \hdashline
 4,5,5 & 14/3\\
 \hdashline
 5,1,1 & 7/3\\
 \hdashline
 5,1,2 & 8/3\\
 \hdashline
 5,1,4 & 10/3\\
 \hdashline
 5,1,5 & 11/3\\
 \hdashline
 5,2,1 & 8/3\\
 \hdashline
 5,2,2 & 3\\
 \hdashline
 5,2,4 & 11/3\\
 \hdashline
 5,2,5 &4\\
 \hdashline
 5,4,1 & 10/3\\
 \hdashline
 5,4,2 & 11/3\\
 \hdashline
 5,4,4 & 13/3\\
 \hdashline
 5,4,5 & 14/3\\
 \hdashline
 5,5,1 &11/3\\
 \hdashline
 5,5,2 & 4\\
 \hdashline
 5,5,4 & 14/3\\
 \hdashline
 5,5,5 & 5\\
 \hdashline
\end{array}"




The sampling distribution of the sample mean "\bar{X}" is



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & \bar{X} & f & f(\bar{X}) & Xf(\bar{X})& X^2f(\bar{X}) \\ \hline
 & 1 & 1 & 1/64 & 1/64 & 3/192\\
 \hdashline
 & 4/3 & 3 & 3/64 & 4/64 & 16/192 \\
 \hdashline
 & 5/3 & 3 & 3/64 & 5/64 & 25/192\\
 \hdashline
 & 2 & 4 & 4/64 & 8/64 & 48/192 \\
 \hdashline
 & 7/3 & 9 & 9/64 & 21/64 & 147/192 \\
 \hdashline
 & 8/3 & 9 & 9/64 & 24/64 & 192/192 \\
 \hdashline
 & 3 & 6 & 6/64 & 18/64 & 162/192 \\
 \hdashline
 & 10/3 & 9 & 9/64 & 30/64 & 300/192 \\
 \hdashline
 & 11/3 & 9 & 9/64 & 33/64 & 363/192 \\
 \hdashline
 & 4 & 4 & 4/64 & 16/64 & 192/192 \\
 \hdashline
 & 13/3 & 3 & 3/64 & 13/64 & 169/192 \\
 \hdashline
 & 14/3 & 3 & 3/64 & 14/64 & 196/192 \\
 \hdashline
 & 5 & 1 & 1/64 & 5/64 & 75/192 \\
 \hdashline
 Sum= & & 64 & 1 & 3 & 59/6 
 
 
 
 
 
 
 
 
 \\
 \hdashline
\end{array}"



e. The mean of the sample means is



"\mu_{ \bar{X}}=E(\bar{X})=3"





"Var(\bar{X})=\sigma^2_{\bar{X}}=E(\bar{X}^2)-(E(\bar{X}))^2"




"=\dfrac{59}{6}-(3)^2=\dfrac{5}{6}""\mu_{\bar{X}}=\mu""\sigma_{\bar{X}}^2=\dfrac{\sigma^2}{n}"


f. Standard deviation of the sampling distribution of sample means:



"\sigma_{\bar{X}}=\sqrt{\sigma_{\bar{X}}^2}=\sqrt{\dfrac{5}{6}}\approx0.91287"